OpenStudy (anonymous):

given the symbol equation to show the formation of iron sulphide: Fe+S=FeS calculate the mass in g of iron sulphide formed when 28g of iron combines with 16g of sulphur

OpenStudy (anonymous):

28/56 = 16/32= 0.5 moles used use that in the equation No.Moles= mass/mr value

OpenStudy (anonymous):

Fe+S=FeS atomic mass of iron = 56g atomic mass of sulphur =32 atomic mass of FeS = 56+32=88g since both quantities are equimolar that is both iron and sulphur are in equal quantity so we can use any to calculate the mass of iron so we can say if 56 g of iron produces 88g of FeS then 28 g will produce = 28 *88/ 56 =2464 /56 = 44g of FeS bingo :)

OpenStudy (anonymous):

thats right @warz

OpenStudy (anonymous):

you are right too just gav a precise ans :D

OpenStudy (anonymous):

yeah....i was lazy :D

OpenStudy (anonymous):

hahahaha i am more lazy i should be preparing for my exam tomorrow just didn't feel like it so ended up here

OpenStudy (anonymous):

haha same here...hi5, i am having an exam..lol but i am already prepared :D

OpenStudy (anonymous):

have*

OpenStudy (anonymous):

:D which subject? i havent studied a chap even but like always i have no tension :)

OpenStudy (anonymous):

thazzzzzzzzzzzzzzzz